A car is involved in a single vehicle accident and the tyre marks were more than 75 metres long. Under road conditions the distance d ( in metres). it takes a car travelling at v m/s to stop is approximated by the equation d= 0.05v^2 + v . The accident occurred in a 100 km/h speed zone. was there a likelihood of the driver exceeding this speed limit ?
Found 2 solutions by Boreal, jorel1380:
Answer by Boreal(15235)

75=0.05v^2+v; multiply by 20 both sides
1500=v^2+20v
v^2+20v-1500=0
(v+50)(v-30)=0
v=30 m/sec only positive root.
That speed is 108,000 m/3600 sec (1 hour)
That is 108 km/h and consistent with the driver's exceeding the speed limit.

Answer by jorel1380(3719)

If d=0.05vē+v, then:
75=0.05vē+v
0.05vē+v-75=0
5vē+100v-7500=0
vē+20v-1500=0
(v+50)(v-30)=0
v=30 or -50
The car was going 30 m/s, or 108 kmh
☺☺☺☺